OpenStudy (anonymous):

using implicit differentiation differentiate tan^3 xy^2+y=x

OpenStudy (dumbcow):

\[\tan^{3} x(y^{2}) +y = x\] is that correct?

OpenStudy (dumbcow):

I will assume it is \[\rightarrow 3y^{2}\tan^{2}x \sec^{2}x + 2y \tan^{3}x \frac{dy}{dx} + \frac{dy}{dx} = 1\] solve for dy/dx \[\frac{dy}{dx} = \frac{1 - 3y^{2} \tan^{2} x \sec^{2}x}{1+2y \tan^{3}x}\]

OpenStudy (anonymous):

3tan^2x sec^2x y^2 + tan^3x (2yy') + y' = 1 tan^3x (2yy') + y' = 1 - 3tan^2x sec^2x y^2 (2tan^3xy + 1) y' = 1 - 3tan^2x sec^2x y^2 => y' = {1 - 3tan^2x sec^2x y^2 / (2tan^3xy + 1)

OpenStudy (anonymous):

@chlorophyll i forgot to introduce brackets at (xy^2+y)